# A Fast Parallel Approach for Neighborhood-based Link Prediction by Disregarding Large Hubs

### v3

@@ 4.2.2. A simple example
Figure 1(a) depicts the standard approach, which considers all second-order neighbors of vertex $1$ for score computation, i.e. vertices $6$ to $12$. This is the process followed by the IBase approach.
Figure 1(b) depicts our DLH approach. It considers only second-order neighbors linked to vertex $1$ by a small hub, with a degree $\leq 8$, for score computation — i.e., vertices $9$ to $12$ which are linked to $1$ through vertices $3$, $4$, and $5$. It is based on the idea, as mentioned above, that low-degree vertices contribute more similarity among their neighbors, while vertices with high degree do not. Note that second-order neighbors of $1$ can still be considered if linked to $1$ by a small hub (e.g., vertex $9$ linked to $1$ through $3$).
Lastly, Figure 1(c) shows the DLH approach, considering only second-order neighbors linked to vertex $1$ by a small hub with a degree $\leq 4$ ($11$ and $12$ linked to $1$ through $4$ and $5$). First order neighbors of $1$ with a degree $>4$ are considered large hubs, and thus their second-order neighbors are disregarded.
@@ 4.2.3. Non-determinism in the result
It’s worth noting that the results obtained by our parallel algorithms (IBase and DLH) are non-deterministic. This is because multiple edges can have matching scores, and the order of edges with the same score may be random due to parallelism.
@@ 4.3. Selecting Suitable Hub Limit for each Similarity Metric
4.4. Our implementation of Disregard Large Hubs (DLH) approach
@@
In the scoring phase (lines 14-33), each vertex is processed in parallel. For each vertex $a$, we scan its second-order neighbors $c$, considering only neighbors of its first order neighbors $b$ which have a degree $|\Gamma_{b}|$ less than or equal to the hub limit $L_{H}$. The scoring of potential edges is done in the scanEdges function (lines 47-53). Here we skip the reverse edges (where $c\leq a$), and calculate a score for each potential second order neighbor $c$ based on the given metric, i.e., for the AA and RA metrics, we apply the scoring formula for each common neighbor $b$, and for the other metrics, we simply count the number of common neighbors. The scores, are accumulated in a collision-free per-thread hashtable $H_{t}$.
@@
4:$a,b,c$: Current vertex, first order, second order neighbor
@@
6:$|\Gamma_{b}|$: Degree of first order neighbor $b$
@@ 5.1.3. Dataset
The graphs used in our experiments are given in Table 1. These are sourced from the SuiteSparse Matrix Collection (Kolodziej et al., 2019). In the graphs, number of vertices vary from $3.07$ to $214$ million, and number of edges vary from $25.4$ million to $3.80$ billion. We ensure edges to be undirected and weighted with a default of $1$.
@@ 5.1.5. Expected precision of Random guess
We now discuss the expected precision of a random guess (instead of using a link prediction algorithm). For an observed graph $G(V,E)$, with $N$ vertices and $M$ edges, there are $N(N-1)/2-M$ possible links, and thus expected the expected precision for predicting $|P|$ edges is $1/{}_{{N(N-1)/2-M}}C_{|P|}$. This is an incredibly small number. For instance, on a graph with $10$ vertices and $100$ edges, correctly predicting all $|P|=10$ has a probability of $3\times 10^{-11}$.
@@ 5.1.6. Measuring Prediction quality
As previously stated, we rely on F1 score, the harmonic mean of precision and recall, to evaluate link prediction performance (Lü et al., 2015). This choice is made due to concerns that AUC may inaccurately favor algorithms ranking many negatives at the bottom (Zhou, 2021; Yang et al., 2015a; Lichtnwalter and Chawla, 2012).
@@ 5.2. Comparative Performance Evaluation
In this section, we compare the performance of the DLH approach, which disregards large hubs, with the IBase approach (which does not). We conduct this comparison for observed graphs $E^{O}$ based on each graph in the dataset, with $10^{-2}|E|$ to $0.1|E|$ unobserved edges. This involves removing $10^{-2}|E|$ to $0.1|E|$ edges, with endpoints chosen uniformly at random (as explained in Section 5.1.4). We predict the same number of edges with both approaches, i.e., $N_{P}=10^{-2}|E|$ or $0.1|E|$. For each observed graph, we predict links using the IBase approach and the DLH approach with a suitable hub limit $L_{H}$ identified in Section 4.3. In Figures 4(a), 4(b), and 4(c), we plot the runtimes, speedups, and F1 scores, respectively, for only the best approach for each graph (considering both F1 score and runtime). In the figures, the labels indicate the abbreviations of the similarity metric used, followed by the value of the hub limit $L_{H}$ parameter setting (e.g., the label $CN32$ stands for the Common Neighbors (CN) similarity metric, with a hub limit $L_{H}$ of $32$, where first order neighbors with a degree greater than $32$ are avoided). It’s worth noting that a hub limit $L_{H}$ of $\infty$ essentially represents the IBase approach. Note that IBase approach crashed on sk-2005 graph with $0.1|E|$ unobserved edges due to out of memory issue, and thus these plots are not shown.
@@
Figure 6. Overall speedup of our approach of Disregarding Large Hubs (DLH) for link prediction, and its phases (obtaining edges with top-k scores per thread, and merging scores from each thread into a common scoreboard), with $10^{-2}|E|$ unobserved edges, with increasing number of threads (in multiples of 2). Increasing the number of threads causes work in merging phase to increase, thus leading to a poor speedup.
@@
Furthermore, we observe that link prediction with the RA metric excels, in terms of both F1 score and runtime, on web graphs and social networks when using the IBase approach. Meanwhile, link prediction with the JC similarity metric outperforms others on road networks and protein k-mer graphs. With the DLH approach and $10^{-2}|E|$ unobserved edges, link prediction with the CN metric (hub limit $L_{H}$ of $32$) excels on web graphs and social networks. For road networks, link prediction with the LHN metric (hub limit $L_{H}$ of $4$) performs the best, and for protein k-mer graphs, link prediction with the JC metric (hub limit $L_{H}$ of $256$) is optimal. However, with $0.1|E|$ unobserved edges, link prediction with the CN metric (hub limit $L_{H}$ of $32$) proves to be the best across all graphs. Figures 8, 9, 10, 11 show the runtimes and F1 scores for link prediction with all similarity measures. Notably, in Figures 10 and 11, DLH approach using the AA and RA metrics performs similarly to the CN metric, but with longer runtimes.
Next, we note that the IBase approach achieves an average F1 score of $1.8\times 10^{-2}$ and $1.1\times 10^{-1}$ when predicting $10^{-2}|E|$ and $0.1|E|$ edges, respectively. In comparison, the DLH approach achieves F1 scores, averaging $3.2\times 10^{-2}$ and $9.8\times 10^{-2}$, respectively. Additionally, we observe that the F1 score tends to be higher on web graphs and social networks but significantly lower on road networks and protein k-mer graphs. This discrepancy is likely due to the average degree of the graphs, as local/neighborhood-based link prediction methods rely on the neighborhood of vertices up to a distance of $2$. Graphs with lower average degree provide less information to such methods for predicting edges. While these F1 scores may seem low (compared to the highest possible F1 score of $1$), it’s important to consider that they are significantly higher than random predictions. Although machine learning-based approaches might achieve higher F1 scores, neighborhood-based similarity measures excel in computational efficiency (both in terms of runtime and space) and interpretability, as mentioned earlier.
@@ 5.4. Strong Scaling
In the final analysis, we evaluate the strong scaling performance of the DLH approach, where we perform link prediction while disregarding large hubs. The assessment involves varying the number of threads from $1$ to $32$ in multiples of $2$ for each input graph. We measure the average time taken to predict $0.1|E|$ links using similarity measures defined in Section 3.1, incorporating the best hub limit $L_{H}$ setting identified in Section 4.3. Figure 6 presents the results, illustrating not only the overall scaling performance but also the scaling of the two phases of each link prediction method: identifying edges with top-$k$ scores in each thread (scoring phase) and combining scores across threads to obtain the global top-$k$ edges (merging phase).
@@ 6. Conclusion
Link prediction aims to anticipate the existence of missing/future connection between nodes based on known interactions and network structure (Arrar et al., 2023). Similarity measures are commonly used for their simplicity, interpretability (Pai et al., 2019; Barbieri et al., 2014), and computational efficiency (Garcia Gasulla and Cortés García, 2014). Further, they are often combined with other approaches (Kumari et al., 2022; AbuOda et al., 2020; Pai et al., 2019), such as ensemble learning (Zhou, 2012).
However, evaluating these algorithms on large networks is crucial for accurate insights (Zhou, 2021; Zhou et al., 2021). Many works in this field focus on small (Guo et al., 2019; Rafiee et al., 2020; Mumin et al., 2022; Papadimitriou et al., 2012; Vega-Oliveros et al., 2021; Saifi et al., 2023; Ferreira et al., 2019; Benhidour et al., 2022) to medium-scale graphs (Yang et al., 2015b; Cui et al., 2016; Kalkan and Hambiralovic, [n. d.]; Mohan et al., 2017; Wang et al., 2019; Bastami et al., 2019; Shin et al., 2012; Garcia Gasulla and Cortés García, 2014), while parallelism becomes essential for larger networks. This technical report addresses both issues and introduces a heuristic for efficient computation. Further, a number of research studies (Gatadi and Swarupa Rani, 2023; Saifi et al., 2023; Benhidour et al., 2022; Mumin et al., 2022; Rafiee et al., 2020; Guo et al., 2019; Yang et al., 2015b; Papadimitriou et al., 2012; Wang et al., 2019) and popular network analysis software (such as NetworKit (Staudt et al., 2016) and igraph (Csardi et al., 2006)), use the baseline approach, which computes similarity scores for all non-connected node pairs and predicts links based on the top-$k$ scores. Despite its simplicity, this approach incurs unnecessary computational costs as many node pairs lack common neighbors, and has a high time complexity.
In this technical report, we first improved upon the baseline approach (IBase). Our parallel IBase approach efficiently finds common neighbors and handles large graphs by tracking top-$k$ edges per-thread and later merging them globally. Additionally, we presented a novel heuristic approach, which disregards large hubs (DLH). It is based on the notion that low-degree nodes contribute significant similarity among neighbors, in contrast to high-degree nodes (interestingly, this is similar to the idea behind the AA and RA similarity metrics). We then experimentally determined suitable hub limits for link prediction with each similarity metric.
Our results indicate that the DLH approach is, on average, over 1622×1622$\times$1622 × and 415×415$\times$415 × faster than the IBase approach with $10^{-2}|E|$ and $0.1|E|$ unobserved edges, respectively. It achieves this speedup while predicting links with an average F1 score that is $80\%$ higher and $13\%$ lower, respectively — meaning similar F1 scores without being too low or high. Notably, on the sk-2005 graph with $0.1|E|$ edges removed, DLH achieves a link prediction rate of $38.1M$ edges/s.
Moreover, RA metric shows superior performance in terms of F1 score and runtime on web graphs and social networks using the IBase approach. On road networks and protein k-mer graphs, JC similarity metric outperforms others with the IBase approach. With the DLH approach and $10^{-2}|E|$ unobserved edges, CN metric (hub limit $L_{H}$ of $32$) excels on web graphs and social networks, LHN metric (hub limit $L_{H}$ of $4$) performs best on road networks, and JC metric (hub limit $L_{H}$ of $256$) is optimal for protein k-mer graphs. However, with $0.1|E|$ unobserved edges, CN metric (hub limit $L_{H}$ of $32$) is the best across all graphs. DLH outperforms IBase, achieving higher F1 scores, particularly on web graphs and social networks.
When predicting $0.1|E|$ edges with the DLH approach, we observed that $63\%$ of the runtime is spent on the scoring phase, and especially higher on social networks, with high average degree. However, a significant portion is also dedicated to the merging phase, suggesting potential optimization opportunities for future work. In the strong scaling analysis, the DLH approach achieves a speedup of 10.4×10.4$\times$10.4 × with 32 threads, indicating a performance increase of 1.6×1.6$\times$1.6 × for every doubling of threads. However, scalability is constrained, particularly due to the merging phase.
In the future, we want to explore optimizing quasi-local and global methods of similarity. While we explore second order neighbors in this report, our techniques can be extended to apply to higher order neighbors.

### v4

@@ 4.2.2. A simple example
Figure 2(a) depicts the standard approach, which considers all second-order neighbors of vertex $1$ for score computation, i.e., vertices $6$ to $12$. This is the process followed by the IBase approach.
Figure 2(b) depicts our DLH approach. It considers only second-order neighbors linked to vertex $1$ by a small hub, with a degree $\leq 8$, for score computation — i.e., vertices $9$ to $12$, which are linked to $1$ through vertices $3$, $4$, and $5$. It is based on the idea, as mentioned above, that low-degree vertices contribute more similarity among their neighbors, while vertices with high degrees do not. Note that second-order neighbors of $1$ can still be considered if linked to $1$ by a small hub (e.g., vertex $9$ linked to $1$ through $3$).
Lastly, Figure 2(c) shows the DLH approach, considering only second-order neighbors linked to vertex $1$ by a small hub with a degree $\leq 4$ ($11$ and $12$ linked to $1$ through $4$ and $5$). First-order neighbors of $1$ with a degree $>4$ are considered large hubs, and thus their second-order neighbors are disregarded.
(a) Our Improved Baseline approach (IBase)
(b) Disregard hubs with degree $>8$ (DLH)
(c) Disregard hubs with degree $>4$ (DLH)
Figure 2. Illustration of our link prediction approach which Disregards Large Hubs (DLH), i.e., $1^{st}$ order neighbors with high degree. Here we focus on the neighborhood of vertex $1$, but the approach applies to each vertex in the graph. In the figure, $1$ is outlined in black, its $1^{st}$ order neighbors in red, its $2^{nd}$ order neighbors in blue, and explored/traversed vertices are shown with a yellow fill. Edge directions indicate traversal, with some $2^{nd}$ order vertices omitted for simplicity (dotted edges). (a) Depicts our Improved Baseline (IBase) approach, which considers all $2^{nd}$ order neighbors of vertex $1$. (b) Presents our DLH approach, which considers only $2^{nd}$ order neighbors linked to $1$ through a small hub (degree $\leq 8$). This pruning reduces runtime and enhances prediction quality. (c) Illustrates the DLH approach, where vertices with degree $>4$ are considered large hubs.
@@ 4.2.3. Non-determinism in the result
It is worth noting that the results obtained by our parallel algorithms (IBase and DLH) are non-deterministic. This is because multiple edges can have matching scores, and the order of edges with the same score may be random due to parallelism.
4.3. Framework of our IHub and DLH approaches
Figure 3 shows the framework of our IBase and DLH approaches. Given an original graph with missing edges, $T$ threads partition the vertices into $T$ sections (using dynamic scheduling). Each thread computes similarity scores between non-connected second-order neighbors for each source vertex, with the option to ignore large hubs (in the DLH approach). Each thread filters the computed scores, retaining only the top $N_{p}$ edges with scores $\geq S_{th}$. The filtered edges are then merged and globally filtered again to keep the top $N_{p}$ edges overall, yielding the predicted edges. These predicted edges can be added to the original graph to obtain an updated graph, which can then be used for further analysis.
Figure 3. Framework diagram of our IBase and DLH approaches for link prediction. Here, $N_{p}$ denotes the maximum number of edges to predict, while $S_{th}$ indicates the threshold score for making a prediction.
@@ 4.4. Selecting Suitable Hub Limit for each Similarity metric
4.5. Our implementation of the Disregard Large Hubs (DLH) approach
@@
In the scoring phase (lines 14-33), each vertex is processed in parallel. For each vertex $a$, we scan its second-order neighbors $c$, considering only neighbors of its first-order neighbors $b$, which have a degree $|\Gamma_{b}|$ less than or equal to the hub limit $L_{H}$. The scoring of potential edges is done in the scanEdges function (lines 47-53). Here, we skip the reverse edges (where $c\leq a$), and calculate a score for each potential second-order neighbor $c$ based on the given metric, i.e., for the AA and RA metrics, we apply the scoring formula for each common neighbor $b$, and for the other metrics, we simply count the number of common neighbors. The scores are accumulated in a collision-free per-thread hashtable $H_{t}$.
@@
4:$a,b,c$: Current vertex, first-order, second-order neighbor
@@
6:$|\Gamma_{b}|$: Degree of first-order neighbor $b$
@@ 5.1.3. Dataset
The graphs used in our experiments are given in Table 1. These are sourced from the SuiteSparse Matrix Collection (Kolodziej et al., 2019). In the graphs, number of vertices vary from $3.07$ to $214$ million, and number of edges vary from $25.4$ million to $3.80$ billion. We ensure that edges are undirected and weighted with a default of $1$.
@@ 5.1.5. Expected precision of Random guess
We now discuss the expected precision of a random guess (instead of using a link prediction algorithm). For an observed graph $G(V,E)$, with $N$ vertices and $M$ edges, there are $N(N-1)/2-M$ possible links, and thus the expected precision for predicting $|P|$ edges is $1/{}_{{N(N-1)/2-M}}C_{|P|}$. This is an incredibly small number. For instance, on a graph with $10$ vertices and $100$ edges, correctly predicting all $|P|=10$ has a probability of $3\times 10^{-11}$.
@@ 5.1.6. Measuring Prediction quality
As previously stated, we rely on the F1 score, the harmonic mean of precision and recall, to evaluate link prediction performance (Lü et al., 2015). This choice is made due to concerns that AUC may inaccurately favor algorithms ranking many negatives at the bottom (Zhou, 2021; Yang et al., 2015a; Lichtnwalter and Chawla, 2012).
@@ 5.2. Comparative Performance Evaluation
In this section, we compare the performance of the DLH approach, which disregards large hubs, with the IBase approach (which does not). We conduct this comparison for observed graphs $E^{O}$ based on each graph in the dataset, with $10^{-2}|E|$ to $0.1|E|$ unobserved edges. This involves removing $10^{-2}|E|$ to $0.1|E|$ edges, with endpoints chosen uniformly at random (as explained in Section 5.1.4). We predict the same number of edges with both approaches, i.e., $N_{P}=10^{-2}|E|$ or $0.1|E|$. For each observed graph, we predict links using the IBase approach and the DLH approach with a suitable hub limit $L_{H}$ identified in Section 4.4. In Figures 5(a), 5(b), and 5(c), we plot the runtimes, speedups, and F1 scores, respectively, for only the best approach for each graph (considering both F1 score and runtime). In the figures, the labels indicate the abbreviations of the similarity metric used, followed by the value of the hub limit $L_{H}$ parameter setting (e.g., the label $CN32$ stands for the Common Neighbors (CN) similarity metric, with a hub limit $L_{H}$ of $32$, where first order neighbors with a degree greater than $32$ are avoided). It’s worth noting that a hub limit $L_{H}$ of $\infty$ essentially represents the IBase approach. Note that the IBase approach crashed on the sk-2005 graph with $0.1|E|$ unobserved edges due to out-of-memory issue, and thus, these plots are not shown.
@@
Figure 7. Overall speedup of our approach of Disregarding Large Hubs (DLH) for link prediction, and its phases (obtaining edges with top-k scores per thread, and merging scores from each thread into a common scoreboard), with $10^{-2}|E|$ unobserved edges, with increasing number of threads (in multiples of 2). Increasing the number of threads causes work in the merging phase to increase, leading to a poor speedup.
@@
Furthermore, we observe that link prediction with the RA metric excels, in terms of both F1 score and runtime, on web graphs and social networks when using the IBase approach. Meanwhile, link prediction with the JC similarity metric outperforms others on road networks and protein k-mer graphs. With the DLH approach and $10^{-2}|E|$ unobserved edges, link prediction with the CN metric (hub limit $L_{H}$ of $32$) excels on web graphs and social networks. For road networks, link prediction with the LHN metric (hub limit $L_{H}$ of $4$) performs the best, and for protein k-mer graphs, link prediction with the JC metric (hub limit $L_{H}$ of $256$) is optimal. However, with $0.1|E|$ unobserved edges, link prediction with the CN metric (hub limit $L_{H}$ of $32$) proves to be the best across all graphs. Figures 9, 10, 11, 12 show the runtimes and F1 scores for link prediction with all similarity measures. Notably, in Figures 11 and 12, the DLH approach using the AA and RA metrics performs similarly to the CN metric, but with longer runtimes.
Next, we note that the IBase approach achieves an average F1 score of $1.8\times 10^{-2}$ and $1.1\times 10^{-1}$ when predicting $10^{-2}|E|$ and $0.1|E|$ edges, respectively. In comparison, the DLH approach achieves F1 scores, averaging $3.2\times 10^{-2}$ and $9.8\times 10^{-2}$, respectively. Additionally, we observe that the F1 score tends to be higher on web graphs and social networks but significantly lower on road networks and protein k-mer graphs. This discrepancy is likely due to the average degree of the graphs, as local/neighborhood-based link prediction methods rely on the neighborhood of vertices up to a distance of $2$. Graphs with lower average degrees provide less information to such methods for predicting edges. While these F1 scores may seem low (compared to the highest possible F1 score of $1$), it’s important to consider that they are significantly higher than random predictions. Although machine learning-based approaches might achieve higher F1 scores, neighborhood-based similarity measures excel in computational efficiency (both in terms of runtime and space) and interpretability, as mentioned earlier.
@@ 5.4. Strong Scaling
In the final analysis, we evaluate the strong scaling performance of our DLH approach, where we perform link prediction while disregarding large hubs, i.e., first order neighbors with high degree. The assessment involves varying the number of threads from $1$ to $32$ in multiples of $2$ for each input graph. We measure the average time taken to predict $0.1|E|$ links using similarity measures defined in Section 3.1, incorporating the best hub limit $L_{H}$ setting identified in Section 4.4. Figure 7 presents the results, illustrating not only the overall scaling performance but also the scaling of the two phases of each link prediction method: identifying edges with top-$k$ scores in each thread (scoring phase) and combining scores across threads to obtain the global top-$k$ edges (merging phase).
@@ 6. Conclusion
Link prediction aims to predict missing or future connections between nodes based on existing interactions and network structure. Similarity measures are widely used for their simplicity, interpretability, and efficiency, and are often combined with methods like ensemble learning. Evaluating these algorithms on large networks is crucial, but many studies focus on small to medium-scale graphs. To address this, we present a parallel approach, IBase, that efficiently finds common neighbors and handles large graphs by tracking top-$k$ edges per thread. We also introduce a novel heuristic that disregards large hubs (DLH). It is based on the notion that low-degree nodes contribute significant similarity among neighbors, in contrast to high-degree nodes (interestingly, this is similar to the idea behind the AA and RA similarity metrics).
Our results show that the DLH approach is, on average, over 1622×1622$\times$1622 × and 415×415$\times$415 × faster than the IBase approach with $10^{-2}|E|$ and $0.1|E|$ unobserved edges, respectively, while maintaining similar F1 scores ($80\%$ higher and $13\%$ lower, respectively). On the sk-2005 graph with $0.1|E|$ edges removed, DLH predicts links at a rate of $38.1M$ edges/s. The RA metric performs best on web graphs and social networks using the IBase approach, while the JC metric excels on road networks and protein k-mer graphs. With DLH and $10^{-2}|E|$, CN (hub limit $L_{H}=32$) works best for web graphs and social networks, LHN ($L_{H}=4$) for road networks, and JC ($L_{H}=256$) for protein k-mer graphs. With $0.1|E|$, CN ($L_{H}=32$) is optimal across all graphs. DLH consistently outperforms IBase, especially on web graphs and social networks. For DLH with $0.1|E|$ edges, $63\%$ of the runtime is spent on scoring, particularly in social networks, with significant time also spent on merging, hinting at optimization opportunities. DLH achieves a 10.4×10.4$\times$10.4 × speedup with $32$ threads, but scalability is limited by the merging phase. Future work may focus on optimizing quasi-local and global similarity methods and extending our techniques to higher-order neighbors.